• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name M. Rvachev

    Log entry time 09:16:35 on December 4,1999

    Entry number 30574

    keyword=Shift summary

    12:35 Hadron Q1 fixed, hall is going to beam permit, waiting for beam from MCC.

    1:15 Hadron Q1 went off during the cycling procedure. Spare interlock and PreReg Fault went on. We took an access a 2:20.

    3:40 Hadron Q3 got a quenched fault when ramping down. We asked for an access
    after calling the tech on call who is coming.
    Q3 fixed.

    Cycled quadrupoles.
    Took optics data.
    Conditions true for the all optics runs:
    energy fast feedback is on;
    beam position fast feedback is off;
    beam spot is appr. 0.5*0.3mm